A 31 g block of ice is cooled to −63 ◦C. It is added to 524 g of water in an 66 g copper calorimeter at a temperature of 26◦C. Find the final temperature. The specific heat of copper is 387 J/kg · ◦C and of ice is 2090 J/kg · ◦C . The latent heat of fusion of water is 3.33 × 105 J/kg and its specific heat is 4186 J/kg · ◦C . Answer in units of ◦C.


1
Expert's answer
2021-05-10T15:28:31-0400

"m_1C_1(0-T_1) - ice\ heating\ to \ 0^oC\\
m1\lambda - ice\ melting\\
m_1C_2(T-0) - melted\ water\ heating\ \\
m_2C_2(T-T_2) - water\ cooling\\
m_3C_3(T-T_3) - copper\ cooling"


"m_1C_1(0-T_1)+m1\lambda+m_1C_2(T-0)+m_2C_2(T-T_2)+m_3C_3(T-T_3)=0\\"

"0.031 \u00b7 2090 \u00b7 63+0.031 \u00b7 3.33 \u00b7 10^5+0.031 \u00b7 4186 \u00b7 T+0.589 \u00b7 2090\u00b7 (T-26)+0.066\u00b7387\u00b7 (T-26) =0"


"1386.328 · T=18265.582\\
T=13.18^oC"

Answer: "T=13.18^oC".